OpenStudy (anonymous):

Divide and Simplify: 4a/7b^2 / 5a^2/8b a. 32/35ab b. 2/35ab c. 5a^3/14b^3 d. 32ab/35

OpenStudy (anonymous):

divide and simplify: \[4a/7b^2 \div 5a^2/8b \]

Parth (parthkohli):

Here's a property: \( \color{Black}{\Rightarrow \Large {a \over b} \div {x \over y} = {a \over b} \times {y \over x}}\) Note how y/x is the reciprocal of x/y and how a/b is unchanged.

OpenStudy (lgbasallote):

Hint: \[\Huge \frac{a}{b} \div \frac{c}{d} = \frac{a}{b} \times \frac{d}{c}\]

Parth (parthkohli):

^ That is not readable to me :P

OpenStudy (lgbasallote):

i have weak 62-year-old eyes okay =_=

OpenStudy (anonymous):

i hate fractions i can never solve them.. can u add in the numbers?

OpenStudy (lgbasallote):

add in? what do you mean by that?

OpenStudy (anonymous):

the numbers i have for my quetion

OpenStudy (lgbasallote):

would you agree that this problem can be rewritten as \[\frac{4a}{7b^2} \div \frac{5a^2}{8b}??\]

OpenStudy (anonymous):

yes thats how it is

Parth (parthkohli):

You can write it as: \(\Huge {4a \over 7b^2} \times {8b \over 5a^2}\) Can't you?

OpenStudy (lgbasallote):

then would you agree that the expression i posted is in the form \[\Huge \frac{w}{x} \div \frac{y}{z}??\] these letters are just mere representations

Parth (parthkohli):

You can also cancel something :)

OpenStudy (anonymous):

im lost again lol... how do u divide it? is it side to side or diagnol

OpenStudy (lgbasallote):

just tell me if you agree that the expression i wrote looks like \(\frac{w}{x} \div \frac{y}{z}\)??

OpenStudy (anonymous):

i guess yea?

OpenStudy (lgbasallote):

then what would be "w" in the expression \(\frac{4a}{7b^2} \div \frac{5a^2}{8b}\)??

OpenStudy (anonymous):

4a?

OpenStudy (lgbasallote):

great! you're right...and what is x?

OpenStudy (anonymous):

7b^2

OpenStudy (lgbasallote):

y?

OpenStudy (lgbasallote):

i mean what's y? not "why"

OpenStudy (anonymous):

5a^2

OpenStudy (lgbasallote):

and z?

OpenStudy (anonymous):

8b

OpenStudy (lgbasallote):

great..so.. w = 4a x = 7b^2 y = 5a^2 z = 8b now remember that i said \[\frac{w}{x} \div \frac{y}{z} = \frac{w}{x} \times \frac{z}{y}\] now..using the letters you assigned...what is \(\large \frac{w}{x} \times \frac{z}{y}\)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

am i cross multiplying?

OpenStudy (lgbasallote):

no no....let's take it from the top...you agreed that \[\frac{4a}{7b^2} \div \frac{5a^2}{8b} \implies \frac{w}{x} \div \frac{y}{z}\] right? you agreed that they are in the same format?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

and i told you that \[\huge \frac{w}{x} \div \frac{y}{z} = \frac{w}{x} \times \frac{z}{y}\] right? i made it bigger so you can see the differences ;) this is actually a rule/formula or whatever you wanna call it lol...

OpenStudy (anonymous):

yes then....?

OpenStudy (lgbasallote):

THEN you shall turn \[\large \frac{4a}{7b^2} \div \frac{5a^2}{8b}\] into the form \[\Large \frac{w}{x} \times \frac{z}{y}\]

OpenStudy (lgbasallote):

note that w = 4a x = 7b^2 y = 5a^2 z = 8b as we stated before

OpenStudy (lgbasallote):

you just have to substitute 4a into w; 7b^2 into x; 5a^2 into y and 8b into z

OpenStudy (anonymous):

would it be answers a or d?

OpenStudy (anonymous):

would it be between those 2?

OpenStudy (anonymous):

is it a?

OpenStudy (anonymous):

nvm i have the answer now but thanks :)

OpenStudy (lgbasallote):

great!!